Proof Boole's inequality
for n events A1,A2,...An.
1
Expert's answer
2020-09-01T18:25:26-0400

Let "B_i=A_i-\cup_{k=1}^{i-1}A_k"

then "B_i\subseteq A_i\ and\ P(B_i)\leq P(A_i)."

Let "i\neq j" , without loss of generality, assume "i>j", then

"B_i\cap B_j=(A_i-\cup_{k=1}^{i-1}A_k)\cap (A_j-\cup_{k=1}^{j-1}A_k)\subseteq\\
\subseteq(A_i-\cup_{k=1}^{i-1}A_k)\cap A_j,"

since "j<i", then "A_j\subseteq \cup_{k=1}^{i-1}A_k", hence "(A_i-\cup_{k=1}^{i-1}A_k)\cap A_j=\empty," and

"B_i\cap B_j=\empty."

Since "B_i\cap B_j=\empty", then, by countable additivity,

"P(\cup_{k=1}^{n}B_k)=\sum_{k=1}^nP(B_k)".

"\cup_{k=1}^{n}B_k=\\
A_1\cup (A_2-A_1)\cup(A_3-\cup_{k=1}^{2}A_k)\cup...\\
...\cup (A_n-\cup_{k=1}^{n-1}A_k)=\\
(A_1\cup A_2)\cup (A_3-\cup_{k=1}^{2}A_k)\cup...\\
...\cup (A_n-\cup_{k=1}^{n-1}A_k)=\\
(A_1\cup A_2\cup A_3)\cup... \cup (A_n-\cup_{k=1}^{n-1}A_k)=...\\
=(A_1\cup A_2\cup...\cup A_{n-1}\cup (A_n-\cup_{k=1}^{n-1}A_k)=\\
\cup_{k=1}^{n}A_k)"

therefore,

"\cup_{k=1}^{n}B_k=\cup_{k=1}^{n}A_k\ and\ \\
P(\cup_{k=1}^{n}A_k)=P(\cup_{k=1}^{n}B_k)=\\
\sum_{k=1}^nP(B_k)\leq \sum_{k=1}^nP(A_k)",

thus,

"P(\cup_{k=1}^{n}A_k)
\leq \sum_{k=1}^nP(A_k)."